Factoring Trinomials by Trial and Error
8x^2-28x+20

Answer by funmath(2873)

Factoring Trinomials by Trial and Error
8x^2-28x+20
First factor the GCF, 4
4(2x^2-7x+5)
The factors of the first term, 2x^2 are:
2x*x
These are the the first term in each paragraph.
The factors of the last term, +5 are the same sign, but you have to add them to get a -7, therefore they're both negative:
-5*-1 and -1*-5
These are the possible second term in each parenthesis:
(2x-5)(x-1)=2x^2-2x-5x+5=2x^2-7x+5 Correct, on the first trial. It's not always that easy.
Therefore,
8x^2-28x+20=4(2x-5)(x-1)
Happy Calculating!!!